OpenStudy (anonymous):

What is the resulting equation when the expression for y in the second equation is substituted into the first equation? 3x + y = 1 y = 6 - 4x -x + 6 = 1 or x + 6 = 1 or 7x + 6 = 1

OpenStudy (anonymous):

You already have part of the solution; substitute the y = 6-4x for the y value in the first equation. Simplify to solve.

OpenStudy (texaschic101):

you can do this...do what Mike said...sub 6 - 4x in for y in the 1st equation, then simplify

OpenStudy (anonymous):

k i got it thx

OpenStudy (texaschic101):

what did you get ?

OpenStudy (anonymous):

-x+6

OpenStudy (texaschic101):

3x + y = 1 3x + (6 - 4x) = 1 3x + 6 - 4x = 1 -x + 6 = 1

OpenStudy (texaschic101):

your right